OpenStudy (anonymous):

I need to find Absolute max and min f(x) = sin2x + 2cosx on [0, π/2]

OpenStudy (tkhunny):

Is that \(sin^{2}(x)\;or\;\sin(2x)\)? Have you considered a 1st Derivative? You WILL have to check the endpoints.

OpenStudy (anonymous):

sin(2x)

OpenStudy (anonymous):

i got cos2x + 2sinx for derivative

OpenStudy (tkhunny):

See how much more clear it is WITH the parentheses? Now, about that derivative... s/b \(f'(x) = 2\cos(2x) - 2\sin(x)\) Something went wrong with both pieces.

OpenStudy (anonymous):

how u got 2cos(2x)?

OpenStudy (tkhunny):

It's the Chain Rule. \(\dfrac{d}{dx}\sin(2x) = \cos(2x)\cdot \dfrac{d}{dx}(2x)\)

OpenStudy (anonymous):

oh thanks

OpenStudy (anonymous):

how do i get the max and min

OpenStudy (anonymous):

do i out 2cos(2x) = 0 and -2sinx = 0 ?

OpenStudy (anonymous):

put*

OpenStudy (tkhunny):

No, that works ONLY with FACTORS, not terms. If it were me, I would convert \(\cos(2x)\) to \(1−2\sin^{2}(x)\) and treat it like a quadratic equation.

OpenStudy (anonymous):

can you please show me all the way? I dont understand much